OpenStudy (anonymous):

HELP PLEASE 3 to the power of 2 over 3 or 3^2/3

OpenStudy (anonymous):

@phi @SolomonZelman @jdoe0001

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{3^2}{3} =\frac{3^2}{3^1}=3^{2-1}=3^1=3 }\)

OpenStudy (anonymous):

MY Answer choices are \[\sqrt[3]{9}\] \[\sqrt[2]{9}\]

OpenStudy (anonymous):

an 2 more here

OpenStudy (solomonzelman):

Ohh \(\normalsize\color{blue}{ 3^{2/3} }\) right ?

OpenStudy (anonymous):

\[\sqrt[3]{27}\] \[\sqrt[2]{27}\]

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3^{2/3} =\sqrt[3]{3^2}=\sqrt[3]{9} }\)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (solomonzelman):

Anytime

OpenStudy (anonymous):

help with one more @SolomonZelman please

OpenStudy (solomonzelman):

Sure...

OpenStudy (anonymous):

\[\sqrt{64}+\sqrt{5}\] \[\sqrt{64}+\sqrt{4}\] Which statement best describes the two expressions? Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational.

OpenStudy (solomonzelman):

What do you think ? I will give you a 2 hint 1 hint, pic

OpenStudy (anonymous):

Both are irrational?

OpenStudy (solomonzelman):

Hint 2: \(\normalsize\color{black}{ \sqrt{64}=8 }\) and \(\normalsize\color{black}{ \sqrt{4}=2,~~~~~~~~~~~~ }\)\(\normalsize\color{black}{ \sqrt{5}=~~\rm{irrational} }\)

OpenStudy (solomonzelman):

What do you think now ?v

OpenStudy (anonymous):

d

OpenStudy (anonymous):

is that correct @SolomonZelman

OpenStudy (anonymous):

hello @SolomonZelman